OpenStudy (anonymous):

Given the function f(x) = The quantity of 4x - 2/3 which of the below expressions is correct? answer choices: f−1(x) = The quantity of 2 - 4x/ 3 f−1(x) = The quantity of 3x - 2 / 4. f−1(x) = The quantity of -4x - 2 / 3. f−1(x) = The quantity of 3x + 2 / 4.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

replace f(x) with y swap x and y then solve for y to get the inverse

OpenStudy (anonymous):

ok what do you mean by swap x and y

jimthompson5910 (jim_thompson5910):

say you had the equation y = 2x + 3 you'd swap x and y to get x = 2y + 3

jimthompson5910 (jim_thompson5910):

then you solve for y in x = 2y + 3 to get the inverse of y = 2x + 3

OpenStudy (anonymous):

oh ok i got C?

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{4x-2}{3}\] \[\Large y = \frac{4x-2}{3}\] \[\Large x = \frac{4y-2}{3}\] Now solve for y

OpenStudy (anonymous):

i dont know the steps

jimthompson5910 (jim_thompson5910):

check out http://regentsprep.org/regents/math/algebra/ae2/lsolveq.htm

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

do you think you could give me an example that is just like this so i can plug in the numbers. thats how i learn best

jimthompson5910 (jim_thompson5910):

Similar Example \[\Large f(x) = \frac{17x-94}{15}\] \[\Large y = \frac{17x-94}{15}\] \[\Large x = \frac{17y-94}{15}\] \[\Large 15x = 17y-94\] \[\Large 15x+94 = 17y\] \[\Large 17y = 15x+94\] \[\Large y = \frac{15x+94}{17}\] \[\Large f^{-1}(x) = \frac{15x+94}{17}\] So the inverse of \(\Large f(x) = \frac{17x-94}{15}\) is \(\Large f^{-1}(x) = \frac{15x+94}{17}\)

OpenStudy (anonymous):

perfect. just what i needed. thanks i got D and i know its correct

jimthompson5910 (jim_thompson5910):

correct